# On schurity of the dihedral group $\boldsymbol{D_{2p}}$

### v3

@@ Abstract.
A finite group $G$ is called a Schur group if every $S$-ring over $G$ is schurian, i.e. associated in a natural way with a subgroup of $\operatorname{Sym}(G)$ that contains all right translations. One of the crucial open questions in the $S$-ring theory is the question on existence of an infinite family of nonabelian Schur groups. We prove that a dihedral group of order $2p$, where $p$ is a Fermat prime or prime of the form $p=4q+1$, where $q$ is also prime, is Schur. Towards this result, we prove nonexistence of a difference set in a cyclic group of order $p\neq 13$ and classify all $S$-rings over some dihedral groups.
Keywords: Schur rings, Schur groups, difference sets.
@@
The author is supported by The Israel Science Foundation (project No. 87792731)
@@ 1. Introduction
A Schur ring or $S$-ring over a finite group $G$ can be defined as a subring of the group ring $\mathbb{Z}G$ that is a free $\mathbb{Z}$-module spanned by a partition of $G$ closed under taking inverse and containing the identity element $e$ of $G$ as a class. The first construction of such ring was proposed by Schur [32]. The general theory of $S$-rings was developed by Wielandt [34]. Schur and Wielandt used $S$-rings to study permutation groups containing regular subgroups. To the present moment, $S$-rings have become a useful tool for studying Cayley graphs, in particular, the Cayley graph isomorphism problem. On the other hand, the theory of $S$-rings continues to develop itself. Concerning the theory of $S$-rings, we refer the readers to the survey [22].
@@
One of the crucial properties of $S$-rings is the schurity property. If $K$ is a subgroup of the symmetric group $\operatorname{Sym}(G)$ containing the group $G_{r}$ of all right translations, then the partition of $G$ into the orbits of the stabilizer $K_{e}$ defines an $S$-ring over $G$. As Wielandt wrote [35], Schur had conjectured that every $S$-ring can be constructed in such way from a suitable permutation group. However, this conjecture was disproved by Wielandt [34]. In honor of this Schur’s fallacy, $S$-rings arising from permutation groups were called schurian, whereas groups whose all $S$-rings are schurian were called Schur [27]. The schurity property for $S$-rings is closely related to the isomorphism poblem for Cayley graphs (see [6, 22]). The problem of determining of all Schur groups was posed by Pöschel [27]. In general, this problem seems hard, in particular, because the number of $S$-rings over a given group is exponential in the order of this group.
@@
The most of the results on Schur groups are concerned with abelian groups. They can be found in [10, 11, 12, 19, 23, 25, 27, 30]. Necessary conditions of schurity for nonabelian groups were obtained in [26]. In [31], it was shown that every nilpotent Schur group belongs to one of the explicitly given infinite families of groups. Nonschurity of some nonabelian groups was checked in [23, 29]. Computer calculations [36] imply that there are nonabelian Schur groups. The largest known nonabelian Schur group has order $63$. It looks natural to ask the following question.
Question.
Does an infinite family of nonabelian Schur groups exist?
The above question seems to be difficult. One reason for this is that the theory of $S$-rings over nonabelian groups is almost undeveloped. Another one is that attempts to find the required family meet some problems which are hard in general, e.g. the problems of existence of a difference set and computation of cyclotomic numbers. The first of them will be discussed further.
@@
In this paper, we are interested in schurity of dihedral groups. Schurity of dihedral $2$-groups was studied in [23]. Namely, in that paper all $S$-rings of rank at most $5$ over dihedral $2$-groups were classified and it was shown that only $S$-rings associated with divisible difference sets can be nonschurian among them. More precisely, in the present paper we are interested in schurity of a dihedral group of order $2p$, where $p$ is prime. In [26], it was proved that such group is not Schur whenever $p\equiv 3\mod~{}4$ and $p>11$. The main result of this paper is the following theorem which is the first noncomputational result stating an existence of nonabelian Schur groups.
@@
Towards the proof of Theorem 1.1, we classify all $S$-rings over a dihedral group of order $2p$, where $p$ is a Fermat prime, i.e. a prime of the form $p=2^{k}+1$, where $k\geq 1$, or $p=rq+1$ for a prime $q$ and $r\in\{2,4\}$ (Theorem 4.1). Recall that primes of the form $12q+$ 1, where $q$ is also prime, are known as safe primes (see, e.g., [28]).
It is not known whether there are infinitely many Fermat primes or primes of the form $rq+1$, where $q$ is also prime and $r\in\{2,4\}$. Only five Fermat primes are known; the largest of them is $65537$. The question whether a number of primes of the form $rq+1$ is infinite is closely related to the famous Dickson conjecture [7] (see also [28, p. 372]). Namely, if the Dickson conjecture is true, then there are infinitely many primes of the form $14q+$ 1. It should be mentioned that the Dickson conjecture is widely believed. For example, the Green-Tao theorem [14] stating that the sequence of prime numbers contains arbitrarily long arithmetic progressions can be considered as a great support to the Dickson conjecutre. Note that the discussed family of primes is infinite modulo the Bateman-Horn conjecture (see [1, 17]) or the generalized Hardy-Littlewood conjecture (see [15]). Computational results on primes $q$ such that $12q+$ 1 or $14q+$ 1 is also prime can be found on the web-page [24].
One of the crucial questions which appears during the studying of schurity of dihedral groups is the question on existence of a difference set in a cyclic group (see [23, 26]). A subset $D$ of the group $G$ is called a difference set if there is a nonnegative integer $\lambda$ such that every nonidentity element $g\in G$ has exactly $\lambda$ representations in the form $g=d_{1}^{-1}d_{2}$, $d_{1},d_{2}\in D$. Difference sets were introduced by Singer [33] and systematic studying of difference sets was started with the papers of Hall [16] and Bruck [5]. For details on difference sets, we refer the readers to [3, Chapter VI]. The keynote ingredient in the proof of Theorem 1.1 is the following theorem.
@@
Note that Theorem 1.2 for a Fermat prime $p$ is very easy and immediately follows from a simple counting argument, whereas the proof of this theorem for $p=4q+1$ requires some technique of working with multipliers of difference sets. The condition $q>3$ in Theorem 1.2 is essential because there is a nontrivial difference set in a cyclic group of order $13$ (see Remark $1$).
We finish the introduction with a brief outline of the paper. Section $2$ contains a necessary background of $S$-rings. In Section $3$, we provide basic definitions and facts concerned with difference sets and prove Theorem 1.2. In Section $4$, we classify all $S$-rings over a dihedral group of order $2p$, where $p$ is a Fermat prime or $p=rq+1$ for a prime $q$ and $r\in\{2,4\}$, and prove Theorem 1.1. Section $5$ contains some further remarks.
The author would like to thank Prof. Ilia Ponomarenko for a comment on Lemma 4.3 and Prof. Gareth Jones for a comment on the Bateman-Horn conjecture. The author is also grateful to Dr. Saveliy Skresanov for a comment on Lemma 5.1.
@@ 2. $S$-rings
In this section, we follow [23] in general. Let $G$ be a finite group and $\mathbb{Z}G$ the integer group ring. The identity element and the set of all non-identity elements of $G$ are denoted by $e$ and $G^{\#}$, respectively. The symmetric group of the set $G$ is denoted by $\operatorname{Sym}(G)$. If $K\leq\operatorname{Sym}(G)$, then the set of all orbits of $K$ on $G$ is denoted by $\operatorname{Orb}(K,G)$. The subgroup of $\operatorname{Sym}(G)$ induced by the right multiplications of $G$ is denoted by $G_{r}$. If $X\subseteq G$, then the element $\sum\limits_{x\in X}{x}$ of the group ring $\mathbb{Z}G$ is denoted by $\underline{X}$. The set $\{x^{-1}:x\in X\}$ is denoted by $X^{-1}$.
@@
The elements of $\mathcal{S}$ are called the basic sets of $\mathcal{A}$. The number of basic sets is called the rank of $\mathcal{A}$ and denoted by $\operatorname{rk}(\mathcal{A})$. Clearly, $\mathbb{Z}G$ is an $S$-ring. The partition $\{\{e\},G\setminus\{e\}\}$ defines the $S$-ring of rank $2$ over $G$. It is easy to check that if $X,Y\in\mathcal{S}(\mathcal{A})$, then $XY\in\mathcal{S}(\mathcal{A})$ whenever $|X|=1$ or $|Y|=1$. The $S$-ring $\mathcal{A}$ is called symmetric if $X=X^{-1}$ for every $X\in\mathcal{S}(\mathcal{A})$.
@@
Let $X,Y\in\mathcal{S}$. If $Z\in\mathcal{S}$, then the number of distinct representations of $z\in Z$ in the form $z=xy$ with $x\in X$ and $y\in Y$ does not depend on the choice of $z\in Z$. Denote this number by $c^{Z}_{XY}$. One can see that $\underline{X}\cdot\underline{Y}=\sum\limits_{Z\in\mathcal{S}(\mathcal{A})}c^{Z}_{XY}\underline{Z}$. Therefore the numbers $c^{Z}_{XY}$ are the structure constants of $\mathcal{A}$ with respect to the basis $\{\underline{X}:\ X\in\mathcal{S}\}$. The $S$-ring $\mathcal{A}$ is called commutative if $c_{XY}^{Z}=c_{YX}^{Z}$ for all $X,Y,Z\in\mathcal{S}(\mathcal{A})$. If $\mathcal{A}$ is symmetric, then $\mathcal{A}$ is commutative. The following equalities hold
@@
for all $X,Y,Z\in\mathcal{S}(\mathcal{A})$ (see [6, Proposition 2.1.12]).
Let $\mathcal{A}$ and $\mathcal{A}^{\prime}$ be $S$-rings over groups $G$ and $G^{\prime}$, respectively. A bijection $f$ from $G$ to $G^{\prime}$ is called a (combinatorial) isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$ if for every $X\in\mathcal{S}(\mathcal{A})$ there is $X^{\prime}\in\mathcal{S}(\mathcal{A}^{\prime})$ such that $\{(x^{f},y^{f})\in G\times G:~{}yx^{-1}\in X\}=\{(x^{\prime},y^{\prime}):~{}y^{\prime}(x^{\prime})^{-1}\in X^{\prime}\}$. If there is an isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$, we say that $\mathcal{A}$ and $\mathcal{A}^{\prime}$ are isomorphic.
A bijection $f\in\operatorname{Sym}(G)$ is defined to be a (combinatorial) automorphism of $\mathcal{A}$ if for all $x,y\in G$, the basic sets containing the elements $yx^{-1}$ and $y^{f}(x^{-1})^{f}$ coincide. The set of all automorphisms of $\mathcal{A}$ forms a group called the automorphism group of $\mathcal{A}$ and denoted by $\operatorname{Aut}(\mathcal{A})$. One can see that $\operatorname{Aut}(\mathcal{A})\geq G_{r}$ and $\mathcal{A}$ is isomorphic to an $S$-ring over a group $H$ if and only if $\operatorname{Aut}(\mathcal{A})$ has a regular subgroup isomorphic to $H$. A group automorphism $\varphi\in\operatorname{Aut}(G)$ is defined to be a Cayley automorphism of $\mathcal{A}$ if $X^{\varphi}=X$ for every $X\in\mathcal{S}(\mathcal{A})$. It is easy to check that $\varphi\in\operatorname{Aut}(\mathcal{A})$.
@@
$V(K,G)=\operatorname{Span}_{\mathbb{Z}}\{\underline{X}:~{}X\in\operatorname{Orb}(K_{e},~{}G)\},$
where $K_{e}$ is a stabilizer of $e$ in $K$, is an $S$-ring over $G$ as it was proved by Schur [32]. An $S$-ring $\mathcal{A}$ over $G$ is called schurian if $\mathcal{A}=V(K,G)$ for some $K\leq\operatorname{Sym}(G)$ with $K\geq G_{r}$. If $\operatorname{rk}(\mathcal{A})=2$, then $\mathcal{A}=V(\operatorname{Sym}(G),G)$. So the $S$-ring of rank $2$ is schurian. Clearly, two isomorphic $S$-rings are schurian or not simultaneously.
Let $K\leq\operatorname{Aut}(G)$. Then $\operatorname{Orb}(K,G)$ forms a partition of $G$ that defines an $S$-ring $\mathcal{A}$ over $G$. In this case, $\mathcal{A}$ is called cyclotomic and denoted by $\operatorname{Cyc}(K,G)$. It is easy to see that $\operatorname{Cyc}(K,G)=V(G_{r}K,G)$ and hence every cyclotomic $S$-ring is schurian. The next lemma follows from the results of [27] (see also [25, Lemma 2.4]).
@@
Every $S$-ring over a group of prime order is cyclotomic and hence schurian.
Let $H$ be a normal $\mathcal{A}$-subgroup of $G$. The $S$-ring $\mathcal{A}$ is called the wreath product of $\mathcal{A}_{H}$ and $\mathcal{A}_{G/H}$ if every basic set $X$ of $\mathcal{A}$ outside $H$ is a union of some $H$-cosets or, equivalently, $H\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus H}$.
@@
[23, Theorem 3.2] The wreath of two $S$-rings is schurian if and only if each of the operands so is.

### v4

# On schurity of dihedral groups

@@ Abstract.
A finite group $G$ is called a Schur group if every $S$-ring over $G$ is schurian, i.e. associated in a natural way with a subgroup of $\operatorname{Sym}(G)$ that contains all right translations. One of the crucial questions in the $S$-ring theory is the question on schurity of nonabelian groups, in particular, on existence of an infinite family of nonabelian Schur groups. In this paper, we study schurity of dihedral groups. We show that any generalized dihedral Schur group is dihedral and obtain necessary conditions of schurity for dihedral groups. Further, we prove that a dihedral group of order $2p$, where $p$ is a Fermat prime or prime of the form $p=4q+1$, where $q$ is also prime, is Schur. Towards this result, we prove nonexistence of a difference set in a cyclic group of order $p\neq 13$ and classify all $S$-rings over some dihedral groups.
Keywords: Schur rings, Schur groups, difference sets, dihedral groups.
@@
The author was supported by the grant of The Natural Science Foundation of Hebei Province (project No. A2023205045)
@@ 1. Introduction
A Schur ring or $S$-ring over a finite group $G$ can be defined as a subring of the group ring $\mathbb{Z}G$ that is a free $\mathbb{Z}$-module spanned by a partition of $G$ closed under taking inverse and containing the identity element $e$ of $G$ as a class. The theory of $S$-rings was initiated by Schur [40] and developed later by Wielandt [42]. Schur and Wielandt used $S$-rings to study permutation groups containing regular subgroups. To the present moment, $S$-rings have become a useful tool for studying permutation groups and Cayley graphs, in particular, the Cayley graph isomorphism problem. On the other hand, the theory of $S$-rings continues to develop itself. Concerning the latter one, we refer the readers to the survey [27].
@@
One of the keynote properties of $S$-rings is schurity. If $K$ is a subgroup of the symmetric group $\operatorname{Sym}(G)$ containing the group $G_{r}$ of all right translations, then the partition of $G$ into the orbits of the stabilizer $K_{e}$ defines an $S$-ring over $G$. As Wielandt wrote [43], Schur had conjectured that every $S$-ring can be constructed in such way from a suitable permutation group. However, this conjecture was disproved by Wielandt [42]. Later, $S$-rings arising from permutation groups were called schurian, whereas groups whose all $S$-rings are schurian were called Schur groups [33]. The schurity property for $S$-rings is closely related to the isomorphism poblem for Cayley graphs (see [7, 27]). The problem of determining of all Schur groups was posed by Pöschel [33]. In general, this problem seems hard. One of the reasons for this is that to verify schurity of a group $G$, it is needed to describe all $S$-rings over $G$, but the number of $S$-rings over $G$ is exponential in the order of $G$.
@@
Most of the results on Schur groups are concerned with the abelian case. They can be found in [14, 15, 16, 24, 28, 30, 33, 37, 39]. In fact, these results imply a complete classification of abelian Schur groups. However, since the general theory of $S$-rings over nonabelian groups is almost undeveloped, there are only few results on schurity of nonabelian groups. Necessary conditions of schurity for nonabelian groups were obtained in [32]. In [38], it was shown that every nilpotent Schur group belongs to one of the explicitly given families of groups. Nonschurity of some nonabelian groups was checked in [28, 36]. There are several examples of nonabelian Schur groups. The largest known nonabelian Schur group of order $63$ was found by computer calculations [44].
@@
It seems reasonable and interesting to study schurity of different classes of nonabelian groups. In the present paper, we deal with the class of dihedral groups. On the one hand, studying schurity of dihedral groups looks as the most natural step after the abelian case towards the general one. On the other hand, the schurity problem for dihedral groups is closely related to another important problem of the algebraic combinatorics, namely, to the problem of existence of a difference set in a cyclic group. This connection will be discussed further. Apparently, there are only two papers [28] and [32] which are concerned with schurity of dihedral groups. In [28], all $S$-rings of rank at most $5$ over dihedral $2$-groups were classified and it was shown that only $S$-rings associated with divisible difference sets can be nonschurian among them, whereas in  [32], it was proved that a dihedral group of order $2p$, where $p$ is a prime, is not Schur whenever $p\equiv 3\mod~{}4$ and $p>11$.
Recall that a nonabelian group is said to be generalized dihedral if it is a semidirect product of an abelian group and a cyclic group of order $2$, where the nontrivial element of the latter one inverses every element of the former one. The first main result of the paper is the theorem below.
Theorem 1.1.
Every generalized dihedral Schur group is dihedral.
The next result of the paper provides restrictions on the order of a dihedral Schur group. The set of prime divisors of an integer $n$ is denoted by $\pi(n)$.
Theorem 1.2.
Let $n\geq 3$ be a positive integer. Suppose that a dihedral group of order $2n$ is Schur. Then $|\pi(n)|\leq 3$. Moreover,
  1. (1)
    if $|\pi(n)|=2$, then $n$ is a product of a prime and an odd prime power or $n=12$,
  2. (2)
    if $|\pi(n)|=3$, then $n$ is a product of three primes.
To prove Theorems 1.1 and 1.2, we construct new families of nonschurian $S$-rings and prove that some known nonschurian $S$-rings are isomorphic to $S$-rings over dihedral groups. Some of the constructions use difference sets in cyclic groups (see, e.g, Proposition 6.5). More specific restrictions on the order of a dihedral Schur group are given in Corollary 6.8. Due to computational results [44], all dihedral groups of order at most $62$ satisfying the conditions from Theorem 1.2 and Corollary 6.8 are Schur except for the groups $D_{52}$ and $D_{60}$ which are not Schur and discussed in detail in Section 9.1.
As it was mentioned before, the order of the largest known nonabelian Schur group is $63$. This leads us to the question on existence of an infinite family of nonabelian Schur groups which was posed for the first time by Olshansky at the conference “Algebra and combinatorics” (Ekaterinburg, 2013). Attempts to find the required family meet some problems which are hard themselves in general, e.g. the problems of existence of a (partial, divisible) difference set, computation of cyclotomic numbers, description of primitive nonschurian $S$-rings etc. One more of the main results of this paper is the following theorem which is the first result stating an existence of a potentially infinite family of nonabelian Schur groups.
@@
Towards the proof of Theorem 1.3, we classify all $S$-rings over a dihedral group of order $2p$, where $p$ is a Fermat prime, i.e. a prime of the form $p=2^{k}+1$, where $k\geq 1$, or $p=rq+1$ for a prime $q$ and $r\in\{2,4\}$ (Theorem 7.1). Recall that primes of the form $12q+$ 1, where $q$ is also prime, are known as safe primes (see, e.g., [35]). One of the keynote cases in the proof of Theorem 7.1 is the case when Cayley graphs corresponding to some basic sets of an $S$-ring are isomorphic to a Paley graph. To handle this case, we use a recent result of Brouwer and Martin [5] on triple intersection numbers of a Paley graph (see the proof of Lemma 7.8).
It is not known whether there are infinitely many Fermat primes or primes of the form $14q+$ 1, where $q$ is also prime. Only five Fermat primes are known; the largest of them is $65537$. The question whether a number of primes of the form $14q+$ 1 is infinite is closely related to the famous and widely-believed Dickson, Bateman-Horn, and generalized Hardy-Littlewood conjectures (see [2, 9, 18, 22, 35]). Namely, if one of these conjectures is true, then there are infinitely many primes of the form $14q+$ 1. Computational results on primes $q$ such that $14q+$ 1 is also prime can be found on the web-page [29].
One of the crucial questions which appears during the studying of schurity of dihedral groups is the question on existence of a difference set in a cyclic group (see [28, 32]). A subset $D$ of a group $G$ is called a difference set if there is a nonnegative integer $\lambda$ such that every nonidentity element $g\in G$ has exactly $\lambda$ representations in the form $g=d_{1}^{-1}d_{2}$, $d_{1},d_{2}\in D$. Difference sets were introduced by Singer [41] and systematic studying of difference sets was started with the papers of Bruck [6] and Hall [19]. For details on difference sets, we refer the readers to [4, Chapter VI]. The keynote ingredient in the proof of Theorem 1.1 is the following theorem.
@@
Note that Theorem 1.2 for a Fermat prime $p$ immediately follows from a simple counting argument, whereas the proof of this theorem for $p=4q+1$ requires some technique of working with multipliers of difference sets. Namely, in the latter case a group $M$ of multipliers of a nontrivial difference set in a cyclic group $C_{p}$ of order $p$ is a subgroup of $\operatorname{Aut}(C_{p})\cong C_{4q}$. We consider all possibilities for the order of $M$. To obtain a contradiction, we use relations between cyclotomic numbers from the classical paper of Dickson [10] and a recent result of Gordon and Schmidt [17] that guarantees an existence of a nontrivial multiplier. The condition $q>3$ in Theorem 1.2 is essential because there is a nontrivial difference set in a cyclic group of order $13$ (see Remark $1$).
We finish the introduction with a brief outline of the paper. Section $2$ contains a necessary background of $S$-rings. In Section $3$, we provide basic definitions and facts concerned with difference sets and prove Theorem 1.4. The proof of Theorem 1.1 is given in Section $4$. Section $5$ contains a construction of a nonschurian $S$-ring over a group of order $2pn$, where $p$ is a prime and $n\geq 3$, which is used in the proof of Theorem 1.2. In Section $6$, we provide several restrictions on the structure of a dihedral Schur group and prove Theorem 1.2. In Section $7$, we classify all $S$-rings over a dihedral group of order $2p$, where $p$ is a Fermat prime or $p=rq+1$ for a prime $q$ and $r\in\{2,4\}$, and prove Theorem 1.3. Section $8$ contains some further remarks, in particular, on $S$-rings from relative difference sets and their connection with $2$-arc transitive dihedrants and necessary conditions of schurity for Frobenius groups. Throughout the paper, cyclic, dihedral, and elementary abelian groups of order $n$ are denoted by $C_{n}$, $D_{n}$, and $E_{n}$, respectively.
The author is very grateful to Dr. Matan Ziv-Av for the help with computer calculations and Prof. Ilia Ponomarenko for the valuable comments on the text.
@@ 2. $S$-rings
In this section, we follow [28] in general. Let $G$ be a finite group and $\mathbb{Z}G$ the integer group ring. The identity element and the set of all non-identity elements of $G$ are denoted by $e$ and $G^{\#}$, respectively. The symmetric group of the set $G$ is denoted by $\operatorname{Sym}(G)$. If $K\leq\operatorname{Sym}(G)$ and $g\in G$, then the set of all orbits of $K$ on $G$ and the one-point stabilizer of $g$ in $K$ are denoted by $\operatorname{Orb}(K,G)$ and $K_{g}$, respectively. For a set $\Delta\subseteq\operatorname{Sym}(G)$ and a section $S=B/A$ of $G$, we set
$\Delta^{S}=\{f^{S}:~{}f\in\Delta,~{}S^{f}=S\},$
where $S^{f}=S$ means that $f$ permutes the $A$-cosets in $B$ and $f^{S}$ denotes the bijection of $S$ induced by $f$.
Given $g\in G$, the permutation of $G$ induced by the right multiplication on $g$ is denoted by $g_{r}$. The subgroup of $\operatorname{Sym}(G)$ induced by the right multiplications of $G$ is denoted by $G_{r}$. The holomorph $G_{r}\rtimes\operatorname{Aut}(G)$ of $G$ is denoted by $\operatorname{Hol}(G)$. If $X\subseteq G$, then the element $\sum\limits_{x\in X}{x}$ of the group ring $\mathbb{Z}G$ is denoted by $\underline{X}$. The set $\{x^{-1}:x\in X\}$ is denoted by $X^{-1}$.
@@
The elements of $\mathcal{S}$ are called the basic sets of $\mathcal{A}$. The number of basic sets is called the rank of $\mathcal{A}$ and denoted by $\operatorname{rk}(\mathcal{A})$. Clearly, $\mathbb{Z}G$ is an $S$-ring. The partition $\{\{e\},G\setminus\{e\}\}$ defines the $S$-ring $\mathcal{T}_{G}$ of rank $2$ over $G$. It is easy to check that if $X,Y\in\mathcal{S}(\mathcal{A})$, then $XY\in\mathcal{S}(\mathcal{A})$ whenever $|X|=1$ or $|Y|=1$. The $S$-ring $\mathcal{A}$ is called symmetric if $X=X^{-1}$ for every $X\in\mathcal{S}(\mathcal{A})$.
@@
Let $X,Y\in\mathcal{S}$. If $Z\in\mathcal{S}$, then the number of distinct representations of $z\in Z$ in the form $z=xy$ with $x\in X$ and $y\in Y$ does not depend on the choice of $z\in Z$. Denote this number by $c^{Z}_{XY}$. One can see that $\underline{X}\cdot\underline{Y}=\sum\limits_{Z\in\mathcal{S}(\mathcal{A})}c^{Z}_{XY}\underline{Z}$. Therefore the numbers $c^{Z}_{XY}$ are the structure constants of $\mathcal{A}$ with respect to the basis $\{\underline{X}:\ X\in\mathcal{S}\}$. The $S$-ring $\mathcal{A}$ is called commutative if $c_{XY}^{Z}=c_{YX}^{Z}$ for all $X,Y,Z\in\mathcal{S}(\mathcal{A})$. If $\mathcal{A}$ is symmetric, then $\mathcal{A}$ is commutative. The equalities
@@
hold for all $X,Y,Z\in\mathcal{S}(\mathcal{A})$ (see [7, Proposition 2.1.12]).
Let $\mathcal{A}$ and $\mathcal{A}^{\prime}$ be $S$-rings over groups $G$ and $G^{\prime}$, respectively. A bijection $\varphi$ from $\mathcal{S}(\mathcal{A})$ to $\mathcal{S}(\mathcal{A}^{\prime})$ is called an algebraic isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$ if $c_{X^{\varphi}Y^{\varphi}}^{Z^{\varphi}}=c_{XY}^{Z}$ for all $X,Y,Z\in\mathcal{S}(\mathcal{A})$. An algebraic isomorphism from $\mathcal{A}$ to itself is called an algebraic automorphism of $\mathcal{A}$. The following statement is known as the first Schur theorem on multipliers (see [42, Theorem 23.9, (a)]).
Lemma 2.1.
Let $\mathcal{A}$ be an $S$-ring over an abelian group $G$. Then the mapping $X\mapsto X^{(m)}=\{x^{m}:~{}x\in X\}$, $X\in\mathcal{S}(\mathcal{A})$, is an algebraic automorphism of $\mathcal{A}$ for every $m\in\mathbb{Z}$ coprime to $|G|$.
A bijection $f$ from $G$ to $G^{\prime}$ is called a (combinatorial) isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$ if for every $X\in\mathcal{S}(\mathcal{A})$ there is $X^{\prime}\in\mathcal{S}(\mathcal{A}^{\prime})$ such that
$\{(x^{f},y^{f})\in G\times G:~{}yx^{-1}\in X\}=\{(x^{\prime},y^{\prime}):~{}y^{\prime}(x^{\prime})^{-1}\in X^{\prime}\}.$
If there is an isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$, we say that $\mathcal{A}$ and $\mathcal{A}^{\prime}$ are isomorphic.
A bijection $f\in\operatorname{Sym}(G)$ is defined to be a (combinatorial) automorphism of $\mathcal{A}$ if for all $x,y\in G$, the basic sets containing the elements $yx^{-1}$ and $y^{f}(x^{-1})^{f}$ coincide. The set of all automorphisms of $\mathcal{A}$ forms a group called the automorphism group of $\mathcal{A}$ and denoted by $\operatorname{Aut}(\mathcal{A})$. One can see that $\operatorname{Aut}(\mathcal{A})\geq G_{r}$ and $\mathcal{A}$ is isomorphic to an $S$-ring over a group $H$ if and only if $\operatorname{Aut}(\mathcal{A})$ has a regular subgroup isomorphic to $H$. The $S$-ring $\mathcal{A}$ is said to be normal if $G_{r}$ is normal in $\operatorname{Aut}(\mathcal{A})$. If $f\in\operatorname{Aut}(\mathcal{A})$, then
$(Xy)^{f}=Xy^{f}$
for all $X\in\mathcal{S}(\mathcal{A})$ and $y\in G$. If $A$ is an $\mathcal{A}$-subgroup of $G$, then the set of all right $A$-cosets is an imprimitivity system of $\operatorname{Aut}(\mathcal{A})$.
A group isomorphism $\alpha$ from $G$ to $G^{\prime}$ is defined to be a Cayley isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$ if $X^{\alpha}\in\mathcal{S}(\mathcal{A}^{\prime})$ for every $X\in\mathcal{S}(\mathcal{A})$. A group automorphism $\alpha\in\operatorname{Aut}(G)$ is defined to be a Cayley automorphism of $\mathcal{A}$ if $X^{\alpha}=X$ for every $X\in\mathcal{S}(\mathcal{A})$. In this case, $\alpha\in\operatorname{Aut}(\mathcal{A})$.
@@
$V(K,G)=\operatorname{Span}_{\mathbb{Z}}\{\underline{X}:~{}X\in\operatorname{Orb}(K_{e},~{}G)\}$
is an $S$-ring over $G$ as it was proved by Schur [40]. An $S$-ring $\mathcal{A}$ over $G$ is called schurian if $\mathcal{A}=V(K,G)$ for some $K\leq\operatorname{Sym}(G)$ with $K\geq G_{r}$. One can verify that $\mathcal{A}$ is schurian if and only if
$\mathcal{A}=V(\operatorname{Aut}(\mathcal{A}),G).$
It is easy to see that $\mathcal{T}_{G}=V(\operatorname{Sym}(G),G)$ and hence $\mathcal{T}_{G}$ is schurian. Clearly, two isomorphic $S$-rings are schurian or not simultaneously. If $\mathcal{A}$ is a schurian $S$-ring and $S$ is an $\mathcal{A}$-section, then
$\mathcal{A}_{S}=V(\operatorname{Aut}(\mathcal{A})^{S},S)$
and hence $\mathcal{A}_{S}$ is also schurian. A finite group $G$ is said to be Schur if every $S$-ring over $G$ is schurian. The above discussion implies the following lemma.
@@
A section of a Schur group is Schur.
Let $K\leq\operatorname{Aut}(G)$. The set $\operatorname{Orb}(K,G)$ forms a partition of $G$ that defines an $S$-ring $\mathcal{A}$ over $G$. In this case, $\mathcal{A}$ is called cyclotomic and denoted by $\operatorname{Cyc}(K,G)$. It is easy to see that $\operatorname{Cyc}(K,G)=V(G_{r}K,G)$ and hence every cyclotomic $S$-ring is schurian. The first statement of the next lemma follows from the results of [33] (see also [30, Lemma 2.4]), whereas the second one from Lemma 2.1.
@@
Let $\mathcal{A}$ be an $S$-ring over a group of prime order. Then
  1. (1)
    $\mathcal{A}$ is cyclotomic and hence schurian;
  2. (2)
    all nontrivial basic sets of $\mathcal{A}$ are of the same size.
Let $\mathcal{A}$ be an $S$-ring over a group $G$ and $K\leq\operatorname{Aut}(G)$ a group of Cayley isomorphisms from $\mathcal{A}$ to itself. Then due to [7, Lemma 2.3.26], the module $\mathcal{A}^{K}=\operatorname{Span}_{\mathbb{Z}}\{\underline{X}^{K}:~{}X\in\mathcal{S}(\mathcal{A})\}$, where
$X^{K}=\bigcup\limits_{\alpha\in K}X^{\alpha},$
is an $S$-ring over $G$.
Let $S=B/A$ be an $\mathcal{A}$-section of $G$. The $S$-ring $\mathcal{A}$ is called the $S$-wreath product or generalized wreath product of $\mathcal{A}_{B}$ and $\mathcal{A}_{G/A}$ if $A\trianglelefteq G$ and every basic set $X$ of $\mathcal{A}$ outside $B$ is a union of some $A$-cosets or, equivalently, $A\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus B}$. In this case, we write $\mathcal{A}=\mathcal{A}_{B}\wr_{S}\mathcal{A}_{G/A}$. The $S$-wreath product is called nontrivial or proper if $A\neq\{e\}$ and $B\neq G$. If $A=B$, then the $S$-wreath product coincides with the wreath product $\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$ of $\mathcal{A}_{A}$ and $\mathcal{A}_{G/A}$.
Given a section $S=B/A$ of a group $G$ such that $A\trianglelefteq G$ and $S$-rings $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ over $B$ and $G/A$, respectively, such that $S$ is both an $\mathcal{A}_{1}$- and an $\mathcal{A}_{2}$-section, and $(\mathcal{A}_{1})_{S}=(\mathcal{A}_{2})_{S}$, there is a unique $S$-ring $\mathcal{A}$ over $G$ that is the $S$-wreath product with $\mathcal{A}_{B}=\mathcal{A}_{1}$ and $\mathcal{A}_{G/A}=\mathcal{A}_{2}$ (see [11]).
Let $A$ and $B$ be subgroups of $G$ such that $G=A\times B$ and $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ $S$-rings over $A$ and $B$, respectively. Then the partition
$\{X_{1}\times X_{2}:~{}X_{1}\in\mathcal{S}(\mathcal{A}_{1}),~{}X_{2}\in\mathcal{S}(\mathcal{A}_{2})\}.$
of $G$ defines the $S$-ring $\mathcal{A}$ over $G$ called the tensor product of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ and denoted by $\mathcal{A}_{1}\otimes\mathcal{A}_{2}$. The tensor product is called nontrivial if $\{e\}<A<G$ and $\{e\}<B<G$.